find the slope if it exists of the line containing the points (-6,1/2) and (-9,1/2)
Answer by Fermat(136)

If you have two points (x1,y1) and (x2,y2), then the slope of the line between these two points is given by
m = (y1-y2)/(x1-x2)
Using the points given,(-6,1/2) and (-9,1/2)
m = (1/2 - 1/2)/(-6 + 9)
m = 0/3
m = 0
=====
Since the slope is zero, then the line is a horizontal line which cuts the y-axis at y = 1/2.
In fact, y=1/2 is the name of this line.
===================================